The band members sold tickets to their concert performance. There were 16 more than twice as many $6 tickets sold as $5 tickets. The total sales were $1643. How many tickets of each price were sold?
Answer by Paul(988)

5 tickets =x
6 tickets = 16+2x

6(16+2x)+5x=1643
96+12x+5x=1643
96+17x=1643
17x=1547
x=91

16+2(91)=198

Hence, for $5 about 91 tickets were sold, and for $6 about 198 tickets were sold.
Paul.